F=-4i+3j
Find the vector opposite f
What does this mean?

Answer by ikleyn(52778)

.
The vector opposite to F = -4i + 3j is the vector G = 4i - 3j.

It means that the sum of vectors F and G is the zero vector:

F + G = 0 = 0*i + 0*j.